OpenStudy (anonymous):

The measure of one vertex angle of a parallelogram is 18 degrees more than twice the measure of the next vertex angle. Find the measures of the angles of the parallelogram.

OpenStudy (zepp):



OpenStudy (zepp):

What is the sum of interior angles of a parallelogram?

OpenStudy (zepp):

 2*180 = 360

OpenStudy (zepp):

Okay your problem states that one of the vertex angle has 18 degrees more than twice the next one. Let the first vertex angle be x and the second one y x = 2y+18 That would be your equation. Now we know that the sum of interior angles is 360 Then x+x+y+y = 360 or 2x + 2y =360 And now you have two equations, solve :)

OpenStudy (anonymous):

if one vertex angle is x, then the remaining angles must be x+18,x,x+18. sum of adjacent angles of a parallelogram is 180. so 2x+18=180 => x=81 so the angles of parallelogram are 81, 99,81,99

OpenStudy (zepp):

@vamgabu, 81 and 99 don't meet requirements.

OpenStudy (anonymous):

if one vertex angle is x, then the remaining angles must be 2x+18,x,2x+18. sum of adjacent angles of a parallelogram is 180. so 3x+18=180 => x=54 so the angles of parallelogram are 54,126,54,126 my mistake previously i'm sorry:)

OpenStudy (anonymous):

So first I solve x=2y+180 then 2x+2y=360

OpenStudy (zepp):

2y + 18 = x * not 180 :P in fact, use the substituion method You'll get 2(2y+18)+2y =360

OpenStudy (anonymous):

oh yes sorry.

OpenStudy (anonymous):

I get 4y+36+2y=360 5y+36=360

OpenStudy (zepp):

That's good :)

OpenStudy (anonymous):

I get 4y+36+2y=360 5y+36=360 324/5y 5y=64.8 is it 5y or 2y

OpenStudy (zepp):

5y + 36 = 360 5y = 324 y = 64.8 Little mistake there.

OpenStudy (anonymous):

y=64.8 yes thats what I meant

OpenStudy (anonymous):

thanks